OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this 14x + 18y = -54 -4x – 14y = 42

OpenStudy (ash2326):

We have \[ 14x+18y=-54\] \[-4x-14y=42\] We have to use elimination method, Let's notice the coefficients of x and y in both the equations. Let's try to eliminate x Let's multiply first equation by 4 and second by 14 and add both , this will eliminate x Multiply first by 4, we get \[56x+72y=-216\] multiply second by 14 we get \[-56x-196y=588\] now let's add the two \[56x+72y-56x-196y=-216+588\] now we get \[\cancel{56x}+72y-\cancel{56x}-196y=-216+588\] we get \[-124y=372\] divide both sides by 124, we get \[-y=3\] so we get \[y=-3\] now let's substitute y=-3 in any of the equations to find x, say first \[ 14x+18y=-54\] y=-3 \[14x+18(-3)=-54\] we get \[14x-54=-54\] add both sides by 54 \[14x=0\] we get x=0 and y=-3

OpenStudy (ash2326):

canjura did you understand?

OpenStudy (anonymous):

Yes Thank you so much

OpenStudy (ash2326):

Are you new to Open Study?